# Can RLHF be More Efficient with Imperfect Reward Models? A Policy Coverage Perspective

### v1

@@ F.2 Some Useful Lemmas
We use $\pi^{k,n}_{\texttt{OFF}}$ to denote the offline policy computed by Alg. 2 called by Alg. 1 at iteration $(k,n)$ for some $\alpha N<n\leq N$. We denote $\mathcal{E}^{k,n}_{\texttt{OFF}}:=\{\pi^{k,n}_{\texttt{OFF}}=\pi^{k,n}\}$ to be the event that Alg. 2 returns $\pi^{k,n}_{\texttt{OFF}}$ as the policy, and use $\mathcal{E}^{k,n}_{w}:=\{\pi^{*}_{r^{w}}=\pi^{k,n}\}$ to denote the event that Alg. 2 pick and return $\pi^{*}_{r^{w}}$. Besides, we use $\neg\mathcal{E}^{k,n}_{\texttt{OFF}}:=\bigcup_{w\in[W]}\mathcal{E}^{k,n}_{w}$ as a short note for the event that Alg. 2 does not return the offline policy $\pi^{k,n}_{\texttt{OFF}}$. Recall the definition $\Delta(w):=J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{*}_{r^{w}})$, and $\Delta_{\min}=\min_{w\in[W]}\Delta(w)$. We will use $w^{*}$ to denote the index of the task achieves $\Delta_{\min}$ (or any of the tasks if multiple maximizers exist). Given the dataset $\mathcal{D}^{k,n}$ we use $\pi^{k,n}_{\text{mix}}:=\frac{1}{|\mathcal{D}^{k,n}|}\sum_{i,j\in\mathcal{D}^{k,n}}\pi^{i,j}$ to be the uniform mixture policy from $\mathcal{D}^{k,n}$.
Then, we decompose the accumulative value gap depending on whether $\mathcal{E}^{k,n}_{\texttt{OFF}}$ is true or not. We use $\mathbb{I}[\mathcal{E}]$ as the indicator function, which takes value 1 if $\mathcal{E}$ happens and otherwise 0. For any $\widetilde{K}\in[K],\widetilde{N}\in[N]$, we have:
@@
$=$ $\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))+\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n})).$ (19)
@@
We first bound the accumulative error when $\mathbb{I}[\mathcal{E}^{k,n}_{\texttt{OFF}}]=1$. On the good events in Lem. 4.1 and Lem. 4.2 (which holds w.p. $1-\delta$), $\mathbb{I}[\mathcal{E}^{k,n}_{\texttt{OFF}}]=1$ implies
${\widehat{V}}(\pi_{\texttt{OFF}};\mathcal{D}^{k,n})\geq\max_{w\in[W]}{\widehat{V}}(\pi^{*}_{r^{w}};\mathcal{D}^{k,n})\geq\max_{w\in[W]}J_{\beta}(\pi^{*}_{r^{w}})-J_{\beta}(\pi_{\texttt{ref}})=J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\Delta_{\min},$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})\leq$ $\Delta_{\min},$
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})\leq$ $c_{2}{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{k,n}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}^{k,n}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}}.$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})\leq$ $\Delta_{\min}\wedge c_{2}{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{k,n}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}^{k,n}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}}$
@@
$\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))=\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\texttt{OFF}}](\Delta_{\min}\wedge\iota^{k,n}).$ (20)
@@
Next, we bound the accumulative error when $\mathbb{I}[\neg\mathcal{E}^{k,n}_{\texttt{OFF}}]=1$. Note that,
$\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))=\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\sum_{\begin{subarray}{c}w\in[W]\\ \Delta(w)>0\end{subarray}}\mathbb{I}[\mathcal{E}^{n,k}_{w}]\Delta(w)$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\iota^{k,n}\leq$ ${\widehat{V}}^{k,n}(\pi^{k,n}_{\texttt{OFF}};\mathcal{D}^{k,n-1}{})$ (Lem. 4.1 and Lem. 4.2)
@@
$\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))=\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\sum_{\begin{subarray}{c}w\in[W]\\ \Delta(w)>0\end{subarray}}{\mathbb{I}}[\mathcal{E}^{n,k}_{w}]\Delta(w)$
@@
$=$ $\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))+\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\texttt{OFF}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))$
@@
See 4.4
Proof.
Under the choice of $T_{0}=\widetilde{\Theta}(\frac{e^{4R}\kappa^{2}(e^{\frac{2R}{\beta}})\mathcal{C}(\Pi)}{\beta^{2}})$, where we introduce a constant $e^{4R}$ that will be used for normalization later. Here $\widetilde{\Theta}(\cdot)$ at most hides poly-log of $T$. By Thm. 3.2, w.p. $1-\delta$, for any $T_{0}\leq t\leq T$,
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}}^{t})=$ $\widetilde{O}(e^{2R}(1+\kappa(e^{\frac{2R}{\beta}})\sum_{\tau=1}^{t}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{\tau})}{\beta t})\sqrt{\frac{1}{t}})$
$=$ $\widetilde{O}(e^{2R}(1+\kappa(e^{\frac{2R}{\beta}})\sum_{\tau=1}^{T_{0}}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{\tau})}{\beta t}+\kappa(e^{\frac{2R}{\beta}})\sum_{\tau=T_{0}+1}^{t}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{\tau})}{\beta t})\sqrt{\frac{1}{t}})$
$=$ $\widetilde{O}(e^{2R}(1+\kappa(e^{\frac{2R}{\beta}})\frac{\sqrt{\mathcal{C}(\Pi)T_{0}}}{\beta T_{0}}+\kappa(e^{\frac{2R}{\beta}})\sum_{\tau=T_{0}+1}^{t}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{\tau})}{\beta t})\sqrt{\frac{1}{t}})$
$=$ $\widetilde{O}(e^{2R}(1+\kappa(e^{\frac{2R}{\beta}})\sum_{\tau=T_{0}+1}^{t}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{\tau})}{\beta t})\sqrt{\frac{1}{t}}).$ (24)
Next, we prove the above RHS only scales with $\sqrt{\frac{1}{t}}$ by induction. First of all, for $t=T_{0}+1$, under Cond. D.2, we have
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}}^{T_{0}})=\widetilde{O}(e^{2R}\sqrt{\frac{1}{t}}).$
For $t>T_{0}$, since $\alpha=0$, we have $\pi^{t+1}=\pi_{\texttt{OFF}}^{t}$. Suppose we already show that for some $t^{\prime}\geq T_{0}+1$, we have $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}}^{T_{0}})=\widetilde{O}(e^{2R}\sqrt{\frac{1}{t}})$ for all $T_{0}\leq t\leq t^{\prime}$. Then, for $t=t^{\prime}+1$, Eq. (24) implies:
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}}^{t^{\prime}+1})=$ $\widetilde{O}(e^{2R}(1+\kappa(e^{\frac{2R}{\beta}})\sum_{\tau=T_{0}+1}^{t^{\prime}+1}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{\tau})}{\beta{(t^{\prime}+1)}})\sqrt{\frac{1}{{t^{\prime}+1}}})$
$=$ $\widetilde{O}(e^{2R}(1+\kappa(e^{\frac{2R}{\beta}})\frac{e^{2R}}{\beta\sqrt{t^{\prime}+1}})\sqrt{\frac{1}{{t^{\prime}+1}}})$ (Induction condition)
$=$ $\widetilde{O}(e^{2R}\sqrt{\frac{1}{{t^{\prime}+1}}}).$ ($t^{\prime}\geq T_{0}$)
Therefore,
$\sum_{t=T_{0}+1}^{T}J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})=$ $\sum_{t=T_{0}}^{T-1}J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}}^{t})$
$=$ $\widetilde{O}(e^{2R}\sqrt{T}).$
Given that $T_{0}$ is fixed and at most depends on $\log T$, the total regret $\sum_{t=1}^{T}J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})$ grows with $\widetilde{O}(\sqrt{T})$ as long as $T>T_{0}$. ∎
@@ Appendix G Connection between Win Rate and Policy Coverage Coefficient
  • Notably, Eq. (25) has a different implication compared with Eq. (26) that we follow in the algorithm design in Sec. 5. More concretely, Eq. (25) suggests we should also disregard those source policies that strongly dominate $\widetilde{\pi}$ when $\widetilde{\pi}$ is close to $\pi^{*}_{r^{*}}$. This makes sense because those source policies may achieve high rewards or win rates by incurring a high KL divergence with $\pi_{\texttt{ref}}$, and therefore, they may not provide good coverage for $\pi^{*}_{r^{*}}$.
@@ Appendix H Useful Lemmas
Appendix I Missing Experiment Details
@@
Additional Experiment Results

### v2

@@ F.2 Some Useful Lemmas
We use $\pi^{k,n}_{\text{Dstl}}$ to denote the offline policy computed by Alg. 2 called by Alg. 1 at iteration $(k,n)$ for some $\alpha N<n\leq N$. We denote $\mathcal{E}^{k,n}_{\text{Dstl}}:=\{\pi^{k,n}_{\text{Dstl}}=\pi^{k,n}\}$ to be the event that Alg. 2 returns $\pi^{k,n}_{\text{Dstl}}$ as the policy, and use $\mathcal{E}^{k,n}_{w}:=\{\pi^{*}_{r^{w}}=\pi^{k,n}\}$ to denote the event that Alg. 2 pick and return $\pi^{*}_{r^{w}}$. Besides, we use $\neg\mathcal{E}^{k,n}_{\text{Dstl}}:=\bigcup_{w\in[W]}\mathcal{E}^{k,n}_{w}$ as a short note for the event that Alg. 2 does not return the offline policy $\pi^{k,n}_{\text{Dstl}}$. Recall the definition $\Delta(w):=J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{*}_{r^{w}})$, and $\Delta_{\min}=\min_{w\in[W]}\Delta(w)$. We will use $w^{*}$ to denote the index of the task achieves $\Delta_{\min}$ (or any of the tasks if multiple maximizers exist). Given the dataset $\mathcal{D}^{k,n}$ we use $\pi^{k,n}_{\text{mix}}:=\frac{1}{|\mathcal{D}^{k,n}|}\sum_{i,j\in\mathcal{D}^{k,n}}\pi^{i,j}$ to be the uniform mixture policy from $\mathcal{D}^{k,n}$.
Then, we decompose the accumulative value gap depending on whether $\mathcal{E}^{k,n}_{\text{Dstl}}$ is true or not. We use $\mathbb{I}[\mathcal{E}]$ as the indicator function, which takes value 1 if $\mathcal{E}$ happens and otherwise 0. For any $\widetilde{K}\in[K],\widetilde{N}\in[N]$, we have:
@@
$=$ $\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))+\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n})).$ (19)
@@
We first bound the accumulative error when $\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}]=1$. On the good events in Lem. 4.1 and Lem. 4.2 (which holds w.p. $1-\delta$), $\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}]=1$ implies
${\widehat{V}}(\pi_{\text{Dstl}};\mathcal{D}^{k,n})\geq\max_{w\in[W]}{\widehat{V}}(\pi^{*}_{r^{w}};\mathcal{D}^{k,n})\geq\max_{w\in[W]}J_{\beta}(\pi^{*}_{r^{w}})-J_{\beta}(\pi_{\texttt{ref}})=J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\Delta_{\min},$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})\leq$ $\Delta_{\min},$
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})\leq$ $c_{2}{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{k,n}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}^{k,n}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}}.$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})\leq$ $\Delta_{\min}\wedge c_{2}{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{k,n}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}^{k,n}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}}$
@@
$\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))=\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}](\Delta_{\min}\wedge\iota^{k,n}).$ (20)
@@
Next, we bound the accumulative error when $\mathbb{I}[\neg\mathcal{E}^{k,n}_{\text{Dstl}}]=1$. Note that,
$\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))=\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\sum_{\begin{subarray}{c}w\in[W]\\ \Delta(w)>0\end{subarray}}\mathbb{I}[\mathcal{E}^{n,k}_{w}]\Delta(w)$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\iota^{k,n}\leq$ ${\widehat{V}}^{k,n}(\pi^{k,n}_{\text{Dstl}};\mathcal{D}^{k,n-1}{})$ (Lem. 4.1 and Lem. 4.2)
@@
$\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))=\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\sum_{\begin{subarray}{c}w\in[W]\\ \Delta(w)>0\end{subarray}}{\mathbb{I}}[\mathcal{E}^{n,k}_{w}]\Delta(w)$
@@
$=$ $\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))+\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\mathbb{I}[\neg\mathcal{E}^{k,n}_{\text{Dstl}}](J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{k,n}))$
@@ Appendix G Connection between Win Rate and Policy Coverage Coefficient
  • Notably, Eq. (24) has a different implication compared with Eq. (25) that we follow in the algorithm design in Sec. 5. More concretely, Eq. (24) suggests we should also disregard those source policies that strongly dominate $\bar{\pi}$ when $\bar{\pi}$ is close to $\pi^{*}_{r^{*}}$. This makes sense because those source policies may achieve high rewards or win rates by incurring a high KL divergence with $\pi_{\texttt{ref}}$, and therefore, they may not provide good coverage for $\pi^{*}_{r^{*}}$.
@@ Appendix H Useful Lemmas
Appendix I Experiment Details and Additional Results
I.1 Details in Experiment Setup
@@
I.2 Additional Experiment Results
Results under Other Choices for $\text{Alg}_{\text{PO}}$ in Alg. 3
In the following, we report the results with two alternative instaniations of $\text{Alg}_{\text{PO}}$: by optimizing the XPO loss (Xie et al., 2024) or the IPO loss (Azar et al., 2024). All the training setups are the same as the experiments where $\text{Alg}_{\text{PO}}$ is DPO, except that we choose a smaller learning rate 1e-5 for $\text{Alg}_{\text{PO}}$ is IPO.
Remark I.1.
Different from DPO and IPO, XPO is an online algorithm itself and in their original design, the pairs of online data are generated by an online exploration policy and another fixed base policy, respectively. However, empirically, Xie et al. (2024) follow the iterative-DPO and utilize the same online learning policy to generate pairs of online data. This exactly aligns with the no transfer baseline we compete with—instantiating $\text{Alg}_{\text{PO}}$ with XPO in Alg. 3 and setting $W=0$, which we refer as iterative-XPO in this paper.
Without Transfer Purely Exploit ROUGE-Lsum Purely Exploit T5-Large
Iter 1 $52.3\pm 1.0$ $50.4\pm 1.6$ $49.9\pm 0.4$
Iter 2 $55.2\pm 1.4$ $52.3\pm 0.3$ $50.1\pm 0.3$
Iter 3 $55.3\pm 1.1$ $51.8\pm 0.5$ $50.3\pm 0.5$
(a) IPO as $\text{Alg}_{\text{PO}}$ in Alg. 3
Without Transfer Purely Exploit ROUGE-Lsum Purely Exploit T5-Large
Iter 1 $52.3\pm 1.1$ $53.4\pm 0.8$ $50.2\pm 0.3$
Iter 2 $51.6\pm 1.3$ $54.7\pm 1.6$ $49.1\pm 1.3$
Iter 3 $52.2\pm 1.6$ $53.8\pm 2.9$ $49.2\pm 1.1$
(b) XPO as $\text{Alg}_{\text{PO}}$ in Alg. 3
Table 2: Similar to Table 1, we report the win rates (%) of the policies trained by empirical TPO (Alg. 3) competed with 3 baselines, presented across 3 columns. Baseline (I): without transfer, i.e., iterative-IPO or iterative-XPO. Baseline (II): purely utilizing ROUGE-LSum (the lowest-quality source task) in transfer learning. Baseline (III): purely utilizing T5-Large (the highest-quality source task) in transfer learning. Results are averaged with 3 random seeds and 95% confidence levels are reported.
Investigation on Source Task Selection